What is the critical t value (one sample t-test) for a non directional hypothesis with an alpha level of .05 and a sample size of 30?
A. t-critical is 1.699
B. t critical is 2.045
C. t critical is 2.042
D. t critical is 2.457
Answer #1

Given that

sample size = n = 30

Level of significance = a-0.05

Here we want to calculate critical t value for one sample t test for a non directional hypothesis ( i.e. for two tailed hypothesis)

Degrees of freedom = n - 1 = 30 - 1 = 29

We get, Critical value = tc = 2.045................... (By using t table ) ............( ANSWER)

OR By using excel command

=TINV(Probability,Degrees of freedom)

=TINV(0.05,29)

= 2.045

tc = 2.045 .............( ANSWER)

Option B )  t critical is 2.045 is correct ............( ANSWER)